Central Systems desires a weighted average cost of capital of 12 percent. The firm has a before-tax cost of debt of 5 percent and a cost of equity of 15.2 percent. What debt-equity ratio is needed for the firm to achieve its targeted weighted average cost of capital if the tax rate is 20%?

a. 0.67

b. 0.56

c. 0.60

d. 0.40

e. 1.78
Answer #1

After-tax cost of debt=5*(1-tax rate)

=5*(1-0.2)=4%

Let debt be $x

Equity be $y

Total=(x+y)

WACC=Respective cost*Respective weight

12=(x/(x+y)*4)+(y/(x+y)*15.2)

12*(x+y)=4x+15.2y

12x+12y=4x+15.2y

x=(15.2-12)y/(12-4)

=0.40 y

Hence debt-equity ratio=debt/equity

=0.40